# Banach algebras associated to twisted étale groupoids: simplicity and pure infiniteness

### v2

@@ 3. Reduced groupoid Banach algebras
where $\tilde{h}_{U}:r^{-1}(d(U))\to r^{-1}(r(U))$ is the homeomorphism given by $\tilde{h}_{U}(\gamma):=d|_{U}^{-1}(r(\gamma))\gamma$ (see [BKM23, Example 2.17]). This implies that $g*f\in\mathcal{B}(\mathcal{G},\mathcal{L})$ and $\|g*f\|_{\infty}\leq\|g\|_{\infty}\cdot\|f\|_{\infty}$. Thus $\|g*f_{n}-g*f\|_{\infty}\to 0$ whenever $\|f_{n}-f\|_{\infty}\to 0$. Since elements in $C_{c}(U,\mathcal{L})$, $U\in\operatorname{Bis}(\mathcal{G})$, span $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ this proves that $\mathcal{B}(\mathcal{G},\mathcal{L})$ is a left $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$-module continuous in the desired sense. Similarly one checks that $\mathcal{B}(\mathcal{G},\mathcal{L})$ is a continuous right $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$-module, and so a bimodule by associativity of the convolution. Since $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ is dense in $\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$ and closed under the convolution, we conclude that $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})*\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$, $\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})*\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})\subseteq\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$. ∎
@@
For any $U\in S(\mathcal{L})\subseteq\operatorname{Bis}(\mathcal{G})$ and $f\in C_{c}(r(U))*\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$ there is a norm one section $b\in C_{c}(U^{*},\mathcal{L})$ such that $|(b*f)|_{X}|=|f\circ s|_{U}^{-1}|$ as functions on $X$ (the modulus can be dropped in the untwisted case).
@@
(4) implies (3) by putting $E^{\mathcal{R}}_{U}(f):=j_{\mathcal{R}}(f)|_{U}$ for $f\in F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ and $U\in\operatorname{Bis}(\mathcal{G})$. (3) implies (2) because $X\in\operatorname{Bis}(\mathcal{G})$. (2$\Rightarrow$(1) is trivial. Assume (1). We need to show (4). For any $U\in S(\mathcal{L})$ and $f\in\mathfrak{C}_{c}(r(U)\mathcal{G},\mathcal{L})$, taking $b$ as in Lemma 3.5 we get
@@
Now assume that (1)(4) hold. To show that $j_{\mathcal{R}}$ is a $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$-bimodule map fix $f\in F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ and $g\in\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$. Pick $\{f_{n}\}_{n=1}^{\infty}\subseteq\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ converging to $f$ in $\|\cdot\|_{\mathcal{R}}$. Then $\{f_{n}\}_{n=1}^{\infty}$ converges to $j_{\mathcal{R}}(f)$ in $\|\cdot\|_{\infty}$ by continuity of $j_{\mathcal{R}}$. Hence also $g*f_{n}\to g*j_{\mathcal{R}}(f)$ in $\|\cdot\|_{\infty}$ by Lemma 3.4. Thus $g*j_{\mathcal{R}}(f)=\lim_{n\to\infty}g*f_{n}=\lim_{n\to\infty}j_{\mathcal{R}}(gf_{n})=j_{\mathcal{R}}(gf)$. Symmetric considerations give $j_{\mathcal{R}}(fg)=j_{\mathcal{R}}(f)*g$. This in particular implies that $\ker j_{\mathcal{R}}$ is an ideal in $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ and that $E^{\mathcal{R}}_{U}$ is a $C_{0}(X)$-module map, for any $U\in\operatorname{Bis}(\mathcal{G})$. For any $S\subseteq\operatorname{Bis}(\mathcal{G})$ covering $\mathcal{G}$, the equality $\ker j_{\mathcal{R}}=\bigcap_{U\in S}\ker E^{\mathcal{R}}_{U}$ is clear as $E^{\mathcal{R}}_{U}$ is the restriction of $j_{\mathcal{R}}$ to $U$. Since $E^{\mathcal{R}}_{U}$ is a continuous $C_{0}(X)$-module map, we also get $\ker E^{\mathcal{R}}_{U}=\{f\in F_{\mathcal{R}}(\mathcal{G},\mathcal{L}):\text{$E^{\mathcal{R}}_{U}(af)=0$ for all $a\in C_{c}(r(U)),\ U\in S$}\}$. This proves the displayed equalities.
Finally assume that $I$ is an ideal in $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ contained in $\ker E_{X}^{\mathcal{R}}$, and let $f\in I$. For every $a\in C_{c}(r(U))$ with $U\in S(\mathcal{L})$, Lemma 3.5 applied to $a*j_{\mathcal{R}}(f)=j_{\mathcal{R}}(af)\in C_{c}(r(U))*\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$ gives $b\in C_{c}(U^{*},\mathcal{L})$ such that $\|E^{\mathcal{R}}_{X}(baf)\|_{\infty}=\|E^{\mathcal{R}}_{U}(af)\|_{\infty}$. Since $baf\in I$ we have $E^{\mathcal{R}}_{U}(af)=0$ and therefore $f\in\ker(j_{\mathcal{R}})$ by the above description of $\ker(j_{\mathcal{R}})$. Hence $I\subseteq\ker(j_{\mathcal{R}})$. ∎
@@
In view of Proposition 3.6 we see that a reduced groupoid Banach algebra of $(\mathcal{G},\mathcal{L})$ is a Banach algebra $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ completion of $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ such that the inclusion $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})\subseteq\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$ extends to an injective contraction $j_{\mathcal{R}}:F_{\mathcal{R}}(\mathcal{G},\mathcal{L})\to\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$ which is isometric on $C_{0}(X)$. Equivalently, $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ is a groupoid Banach algebra such that, fixing any $S\subseteq\operatorname{Bis}(\mathcal{G})$ covering $\mathcal{G}$, every $f\in F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ is uniquely determined by elements $\{E^{\mathcal{R}}_{U}(f)\}_{U\in S}$ where $E^{\mathcal{R}}_{U}:F_{\mathcal{R}}(\mathcal{G},\mathcal{L})\to\mathcal{B}(U,\mathcal{L})$ is a contraction that extends $E_{U}$, $U\in S$. In fact, by the displayed equality in Proposition 3.6, elements in $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ are determined by the restricted maps $E^{\mathcal{R}}_{U}:C_{0}(r(U))F_{\mathcal{R}}(\mathcal{G},\mathcal{L})\to\mathcal{B}_{0}(U,\mathcal{L})$, $U\in S$. These restricted maps take values in $C_{0}(U,\mathcal{L})$ when $\mathcal{G}$ is Hausdorff. In the context of Fell bundles over inverse semigroups such maps were considered in [KM20, Proposition 2.18]. If $S\subseteq S(\mathcal{L})$ then fixing for every $U\in S$ a unitary section that trivializes the bundle $\mathcal{L}|_{U}$ and using the homeomorphism $r:U\to r(U)$ we get an isomorphism $\Phi_{U}:\mathcal{B}_{0}(U,\mathcal{L})\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\mathcal{B}_{0}(r(U))$ that restricts to $C_{0}(U,\mathcal{L})\cong C_{0}(r(U))$. Thus if in addition $\mathcal{G}$ is Hausdorff we get maps
@@
If $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ is a groupoid Banach *-algebra meaning that the involution on $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ is isometric in $\|\cdot\|_{\mathcal{R}}$, then the map $j_{\mathcal{R}}$ is *-preserving and both the ideal $\ker j_{\mathcal{R}}$ and the quotient reduced algebra $F_{\mathcal{R},\operatorname{r}}(\mathcal{G},\mathcal{L})=F_{\mathcal{R}}(\mathcal{G},\mathcal{L})/\ker j_{\mathcal{R}}$ are Banach *-algebras.
@@
The following lemma generalizes [BK24, Lemma 4.10]. It allows us to produce (reduced) groupoid Banach algebras and -Banach algebras from other (reduced) groupoid Banach algebras.
@@
Inspired by [BK24, Definition 4.12] we now generalize $L^{p}$-groupoid operator algebras to $L^{P}$-algebras where $P\subseteq[1,\infty]$ is a set of Hölder exponents.
@@
By Lemma 3.12 it suffices to consider the case when $P=\{p\}$ is a singleton (in particular, $j_{P}^{\operatorname{r}}=\prod_{p\in P}j_{p}^{\operatorname{r}}\circ\pi$, where $\pi:F^{P}_{\operatorname{r}}(\mathcal{G},\mathcal{L})\to\prod_{p\in P}F^{p}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$ is given by $\pi(a):=\prod_{p\in P}a$, for $a\in\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$). Moreover, since involution on $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ gives $F_{\infty}({\mathcal{G},\mathcal{L}},{\operatorname{r}})\stackrel{{\scriptstyle\text{anti}}}{{\cong}}F_{1}({\mathcal{G},\mathcal{L}},{\operatorname{r}})$ we may assume that $p<\infty$. Then the claim, using a different picture of the twist, in essence follows from [HO23, Propositions 6.21, 6.24]. For the sake of completeness and further reference we give a complete proof based on the argument behind [BFPR21, Proposition 2.8]. Namely, we identify $F^{p}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$ with the corresponding subalgebra of $B(\ell^{p}(\mathcal{G},\mathcal{L}))$ using the extension of the map $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})\to B(\ell^{p}(\mathcal{G},\mathcal{L}))$ given by (2.3). For each $\gamma\in\mathcal{G}$ choose a norm one element $1_{\gamma}\in L_{\gamma}$ and treat it as a section of $\mathcal{L}$ which is zero at $\eta\neq\gamma$. Then $\{1_{\gamma}\}_{\gamma\in\mathcal{G}}$ is a Schauder basis for $\ell^{p}(\mathcal{G},\mathcal{L})$ and for any $f\in\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$ we have $\|f\|_{\infty}=\sup_{\gamma\in\mathcal{G}}|(f1_{d(\gamma)})(\gamma)|\leq\sup_{\gamma\in\mathcal{G}}\|f1_{d(\gamma)}\|_{p}\leq\|f\|_{L^{p},\operatorname{r}}$. Thus we have a contractive linear map $j_{p}^{\operatorname{r}}:F^{p}_{\operatorname{r}}(\mathcal{G},\mathcal{L})\to\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$ extending the inclusion $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})\subseteq\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$. Clearly for every $\gamma\in\mathcal{G}$ and any $f\in\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ we have
@@
By continuity these relations hold for any $f\in F^{p}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$. The second formula in 3.8 implies that $j_{p}^{\operatorname{r}}$ is injective on $F^{p}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$ and hence $F^{p}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$ is a reduced groupoid Banach algebra.
Since yields the isometry $F^{p}_{\operatorname{r}}(\mathcal{G},\mathcal{L})\stackrel{{\scriptstyle\text{anti}}}{{\cong}}F^{q}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$, where $11/p+1/q=$ 1, we get $j_{p}^{\operatorname{r}}(f)^{*}=j_{q}^{\operatorname{r}}(f^{*})$, $f\in F_{p}({\mathcal{G},\mathcal{L}},{r})$. Now we prove (3.7), with $P=\{p\}$. Fix $f,g\in F^{p}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$ and take nets $(f_{i}),(g_{j})$ in $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ converging to $f,g$ in the norm of $F^{p}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$. Using Hölders inequality, (3.8) and the isometry $F^{p}_{\operatorname{r}}(\mathcal{G},\mathcal{L})\stackrel{{\scriptstyle\text{anti}}}{{\cong}}F^{q}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$, for any $\gamma\in\mathcal{G}$ we get
@@
Similarly we get $\sum_{r(\eta)=r(\gamma)}|j_{p}^{\operatorname{r}}(f)(\eta)\cdot j_{p}^{\operatorname{r}}(g_{j}-g)(\eta^{-1}\gamma)|\leq\|f\|_{L^{p},\operatorname{r}}\|g_{j}-g\|_{L^{p},\operatorname{r}}$. Altogether, since $j_{p}^{\operatorname{r}}(f_{i}\cdot g_{j})=f_{i}*g_{j}$, this implies that
@@
If $P=P^{*}$, then $F^{P}(\mathcal{G},\mathcal{L})$ and $F^{P}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$ are Banach *-algebras and $\Lambda_{P}$ is a *-homomorphism, cf. [BKM23, Theorem 5.6(4)]. For $P=\{p,q\}$ with $11/p+1/q=$ 1, the Banach *-algebras $F^{P}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$ are sometimes called symmetrized pseudofunction algebras. They were studied in [AO22] and in the group case in [Elk24], [LY17],[Phi19]. When $P\subseteq[1,\infty]$ contains $\{1,\infty\}$, then $F^{P}(\mathcal{G},\mathcal{L})=F^{P}_{\operatorname{r}}(\mathcal{G},\mathcal{L})=F_{I}(\mathcal{G},\mathcal{L})$, by (2.2), (2.4). In particular, $F_{I}(\mathcal{G},\mathcal{L})$ is a reduced groupoid Banach algebra. In fact, we always have $F^{P}(\mathcal{G},\mathcal{L})=F^{P}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$ for any $P\subseteq\{1,\infty\}$, and if $\mathcal{G}$ is second countable and amenable, then $F^{P}(\mathcal{G},\mathcal{L})=F^{P}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$ for every $P\subseteq[1,\infty]$, see the last part of Example 2.7.

### v3

@@ 3. Reduced groupoid Banach algebras
where $\tilde{h}_{U}:r^{-1}(d(U))\to r^{-1}(r(U))$ is the homeomorphism given by $\tilde{h}_{U}(\gamma):=d|_{U}^{-1}(r(\gamma))\gamma$ (see [BKM25, Example 2.19]). This implies that $g*f\in\mathcal{B}(\mathcal{G},\mathcal{L})$ and $\|g*f\|_{\infty}\leq\|g\|_{\infty}\cdot\|f\|_{\infty}$. Thus $\|g*f_{n}-g*f\|_{\infty}\to 0$ whenever $\|f_{n}-f\|_{\infty}\to 0$. Since elements in $C_{c}(U,\mathcal{L})$, $U\in\operatorname{Bis}(\mathcal{G})$, span $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$, this proves that $\mathcal{B}(\mathcal{G},\mathcal{L})$ is a left $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$-module continuous in the desired sense. Similarly, one checks that $\mathcal{B}(\mathcal{G},\mathcal{L})$ is a continuous right $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$-module, and so a bimodule by associativity of convolution. Since $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ is dense in $\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$ and closed under convolution, we conclude that $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})*\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$, $\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})*\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})\subseteq\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$. ∎
@@
For any $U\in S(\mathcal{L})\subseteq\operatorname{Bis}(\mathcal{G})$ and $f\in C_{c}(r(U))*\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$, there is a norm one section $b\in C_{c}(U^{*},\mathcal{L})$ such that $|(b*f)|_{X}|=|f\circ s|_{U}^{-1}|$ as functions on $X$ (the modulus can be dropped in the untwisted case).
@@
[…]
@@
Now assume that (1)(4) hold. To show that $j_{\mathcal{R}}$ is a $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$-bimodule map fix $f\in F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ and $g\in\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$. Choose a sequence $(f_{n})_{n=1}^{\infty}$ in $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ converging to $f$ in $\|\cdot\|_{\mathcal{R}}$. Then $(f_{n})_{n=1}^{\infty}$ converges to $j_{\mathcal{R}}(f)$ in $\|\cdot\|_{\infty}$ by continuity of $j_{\mathcal{R}}$. Hence also $g*f_{n}\to g*j_{\mathcal{R}}(f)$ in $\|\cdot\|_{\infty}$ by Lemma 3.4. Thus $g*j_{\mathcal{R}}(f)=\lim_{n\to\infty}g*f_{n}=\lim_{n\to\infty}j_{\mathcal{R}}(gf_{n})=j_{\mathcal{R}}(gf)$. Symmetric considerations give $j_{\mathcal{R}}(fg)=j_{\mathcal{R}}(f)*g$. This in particular implies that $\ker j_{\mathcal{R}}$ is an ideal in $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ and that $E^{\mathcal{R}}_{U}$ is a $C_{0}(X)$-module map, for any $U\in\operatorname{Bis}(\mathcal{G})$. For any $S\subseteq\operatorname{Bis}(\mathcal{G})$ covering $\mathcal{G}$, the equality $\ker j_{\mathcal{R}}=\bigcap_{U\in S}\ker E^{\mathcal{R}}_{U}$ is clear as $E^{\mathcal{R}}_{U}$ is the restriction of $j_{\mathcal{R}}$ to $U$. Since $E^{\mathcal{R}}_{U}$ is a continuous $C_{0}(X)$-module map, we also get $\ker E^{\mathcal{R}}_{U}=\{f\in F_{\mathcal{R}}(\mathcal{G},\mathcal{L}):\text{$E^{\mathcal{R}}_{U}(af)=0$ for all $a\in C_{c}(r(U)),\ U\in S$}\}$. This proves the displayed equalities.
Finally, assume that $I$ is an ideal in $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ contained in $\ker E_{X}^{\mathcal{R}}$, and let $f\in I$. For every $a\in C_{c}(r(U))$ with $U\in S(\mathcal{L})$, Lemma 3.5 applied to $a*j_{\mathcal{R}}(f)=j_{\mathcal{R}}(af)\in C_{c}(r(U))*\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$ gives $b\in C_{c}(U^{*},\mathcal{L})$ such that $\|E^{\mathcal{R}}_{X}(baf)\|_{\infty}=\|E^{\mathcal{R}}_{U}(af)\|_{\infty}$. Since $baf\in I$, we have $E^{\mathcal{R}}_{U}(af)=0$ and therefore $f\in\ker(j_{\mathcal{R}})$ by the above description of $\ker(j_{\mathcal{R}})$. Hence $I\subseteq\ker(j_{\mathcal{R}})$. ∎
@@
In view of Proposition 3.6, we see that a reduced groupoid Banach algebra of $(\mathcal{G},\mathcal{L})$ is a Banach algebra $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ completion of $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ such that the inclusion $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})\subseteq\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$ extends to an injective contraction $j_{\mathcal{R}}:F_{\mathcal{R}}(\mathcal{G},\mathcal{L})\to\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$ which is isometric on $C_{0}(X)$. Equivalently, $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ is a groupoid Banach algebra such that, fixing any $S\subseteq\operatorname{Bis}(\mathcal{G})$ covering $\mathcal{G}$, every $f\in F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ is uniquely determined by elements $\{E^{\mathcal{R}}_{U}(f)\}_{U\in S}$ where $E^{\mathcal{R}}_{U}:F_{\mathcal{R}}(\mathcal{G},\mathcal{L})\to\mathcal{B}(U,\mathcal{L})$ is a contraction that extends $E_{U}$, $U\in S$. In fact, by the displayed equality in Proposition 3.6, elements in $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ are determined by the restricted maps $E^{\mathcal{R}}_{U}:C_{0}(r(U))F_{\mathcal{R}}(\mathcal{G},\mathcal{L})\to\mathcal{B}_{0}(U,\mathcal{L})$, $U\in S$. These restricted maps take values in $C_{0}(U,\mathcal{L})$ when $\mathcal{G}$ is Hausdorff. In the context of Fell bundles over inverse semigroups, such maps were considered in [KM20, Proposition 2.18]. Assume that $S\subseteq S(\mathcal{L})$ and for each $U\in S$ fix a unitary section that trivializes the bundle $\mathcal{L}|_{U}$. Then using the homeomorphism $r:U\to r(U)$, we get an isomorphism $\Phi_{U}:\mathcal{B}_{0}(U,\mathcal{L})\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\mathcal{B}_{0}(r(U))$ that restricts to $C_{0}(U,\mathcal{L})\cong C_{0}(r(U))$. Thus if in addition $\mathcal{G}$ is Hausdorff, we get maps
@@
If $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ is a groupoid Banach *-algebra, meaning that the involution on $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ is isometric in $\|\cdot\|_{\mathcal{R}}$, then the map $j_{\mathcal{R}}$ is *-preserving and both the ideal $\ker j_{\mathcal{R}}$ and the quotient reduced algebra $F_{\mathcal{R},\operatorname{r}}(\mathcal{G},\mathcal{L})=F_{\mathcal{R}}(\mathcal{G},\mathcal{L})/\ker j_{\mathcal{R}}$ are Banach *-algebras.
@@
The following lemma generalizes [BK24, Lemma 4.10]. It allows us to produce (reduced) groupoid Banach algebras and Banach-* algebras from other (reduced) groupoid Banach algebras.
@@
Inspired by [BK24, Definition 4.12], we now generalize $L^{p}$-groupoid operator algebras to $L^{P}$-algebras where $P\subseteq[1,\infty]$ is a set of Hölder exponents.
@@
By Lemma 3.12, it suffices to consider the case when $P=\{p\}$ is a singleton (in particular, $j_{P}^{\operatorname{r}}=\prod_{p\in P}j_{p}^{\operatorname{r}}\circ\pi$, where $\pi:F^{P}_{\operatorname{r}}(\mathcal{G},\mathcal{L})\to\prod_{p\in P}F^{p}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$ is given by $\pi(a):=\prod_{p\in P}a$, for $a\in\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$). Moreover, since involution on $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ gives $F_{\infty}({\mathcal{G},\mathcal{L}},{\operatorname{r}})\stackrel{{\scriptstyle\text{anti}}}{{\cong}}F_{1}({\mathcal{G},\mathcal{L}},{\operatorname{r}})$ we may assume that $p<\infty$. Then the claim, using a different picture of the twist, in essence follows from [HO23, Propositions 6.21, 6.24]. For the sake of completeness and further reference, we give a complete proof based on the argument behind [BFPR21, Proposition 2.8]. Namely, we identify $F^{p}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$ with the corresponding subalgebra of $B(\ell^{p}(\mathcal{G},\mathcal{L}))$ using the extension of the map $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})\to B(\ell^{p}(\mathcal{G},\mathcal{L}))$ given by (2.2). For each $\gamma\in\mathcal{G}$, choose a norm one element $1_{\gamma}\in L_{\gamma}$ and treat it as a section of $\mathcal{L}$ which is zero at $\eta\neq\gamma$. Then $\{1_{\gamma}\}_{\gamma\in\mathcal{G}}$ is a Schauder basis for $\ell^{p}(\mathcal{G},\mathcal{L})$ and for any $f\in\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$ we have $\|f\|_{\infty}=\sup_{\gamma\in\mathcal{G}}|(f1_{d(\gamma)})(\gamma)|\leq\sup_{\gamma\in\mathcal{G}}\|f1_{d(\gamma)}\|_{p}\leq\|f\|_{L^{p},\operatorname{r}}$. Thus we have a contractive linear map $j_{p}^{\operatorname{r}}:F^{p}_{\operatorname{r}}(\mathcal{G},\mathcal{L})\to\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$ extending the inclusion $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})\subseteq\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$. Clearly, for every $\gamma\in\mathcal{G}$ and any $f\in\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$, we have
@@
By continuity, these relations hold for any $f\in F^{p}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$. The second formula in 3.7 implies that $j_{p}^{\operatorname{r}}$ is injective on $F^{p}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$ and hence $F^{p}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$ is a reduced groupoid Banach algebra.
Since yields the isometry $F^{p}_{\operatorname{r}}(\mathcal{G},\mathcal{L})\stackrel{{\scriptstyle\text{anti}}}{{\cong}}F^{q}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$, where $1/p+1/q=1$, we get $j_{p}^{\operatorname{r}}(f)^{*}=j_{q}^{\operatorname{r}}(f^{*})$, $f\in F_{p}({\mathcal{G},\mathcal{L}},{r})$. Now we prove (3.6), with $P=\{p\}$. Fix $f,g\in F^{p}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$ and take nets $(f_{i}),(g_{j})$ in $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ converging to $f,g$ in the norm of $F^{p}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$. Using Hölders inequality, (3.7), and the isometry $F^{p}_{\operatorname{r}}(\mathcal{G},\mathcal{L})\stackrel{{\scriptstyle\text{anti}}}{{\cong}}F^{q}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$, for any $\gamma\in\mathcal{G}$, we get
@@
Similarly, we get $\sum_{r(\eta)=r(\gamma)}|j_{p}^{\operatorname{r}}(f)(\eta)\cdot j_{p}^{\operatorname{r}}(g_{j}-g)(\eta^{-1}\gamma)|\leq\|f\|_{L^{p},\operatorname{r}}\|g_{j}-g\|_{L^{p},\operatorname{r}}$. Altogether, since $j_{p}^{\operatorname{r}}(f_{i}\cdot g_{j})=f_{i}*g_{j}$, this implies that
@@
If $P=P^{*}$, then $F^{P}(\mathcal{G},\mathcal{L})$ and $F^{P}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$ are Banach *-algebras and $\Lambda_{P}$ is a *-homomorphism, cf. [BKM25, Theorem 5.13(7)]. For $P=\{p,q\}$ with $1/p+1/q=1$, the Banach *-algebras $F^{P}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$ are sometimes called symmetrized pseudo-function algebras. They were studied in [AO22] and in the group case in [Elk25], [LY17], [Phi19]. When $P\subseteq[1,\infty]$ contains $\{1,\infty\}$, then $F^{P}(\mathcal{G},\mathcal{L})=F^{P}_{\operatorname{r}}(\mathcal{G},\mathcal{L})=F_{I}(\mathcal{G},\mathcal{L})$, by (2.1) and (2.3). In particular, $F_{I}(\mathcal{G},\mathcal{L})$ is a reduced groupoid Banach algebra. In fact, we always have $F^{P}(\mathcal{G},\mathcal{L})=F^{P}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$ for any $P\subseteq\{1,\infty\}$, and if $\mathcal{G}$ is second countable and amenable, then $F^{P}(\mathcal{G},\mathcal{L})=F^{P}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$ for every $P\subseteq[1,\infty]$, see the last part of Example 2.7.